OpenStudy (anonymous):

How do I find the inverse of f(x) when the equation is: f(x)=(x+3)^2, x<-3

OpenStudy (anonymous):

The answer is supposed to be f-1(x)= -√x - 3, but why?

sam (.sam.):

y=(x+3)^{2} \[x=\sqrt{y}-3\] \[f^{-1}x=\sqrt{x}-3\]

OpenStudy (anonymous):

how is y square rooted instead of x? o.O

OpenStudy (apoorvk):

inverse of a function is all about writing y in terms of x.  now interchange x and y to get get the inverse of y.( i dont think there would be a '-' before sqrt(y) in the answer you posted)